# Making Sense of Gravitational Thermodynamics

### v1

@@ Abstract
The use of statistical methods to model gravitational systems is crucial to physics practice, but the extent to which thermodynamics and statistical mechanics genuinely apply to these systems is a contentious issue. This paper provides new conceptual foundations for gravitational thermodynamics by reconsidering the nature of key concepts like equilibrium and advancing a novel way of understanding thermodynamics. The challenges arise from the peculiar characteristics of the gravitational potential, leading to non-extensive energy and entropy, negative heat capacity, and a lack of standard equilibrium. Hence it has been claimed that only non-equilibrium statistical mechanics is warranted in this domain, whereas thermodynamics is inapplicable. We argue instead that equilibrium statistical mechanics applies to self-gravitating systems at the relevant scale, as they display equilibrium in the form of metastable quasi-equilibrium states. We then develop a minimal framework for thermodynamics that can be applied to these systems and beyond. Thermodynamics applies in the sense that we can devise macroscopic descriptions and explanations of the behaviour of these systems in terms of coarse-grained quantities within equilibrium statistical mechanics.
@@ 1 Introduction
Second, although full-blown phenomenological thermodynamics is unsuitable in this domain, we develop what we call a ‘minimal framework for thermodynamics’ and show how a notion of thermodynamics applies to SGS. In fact, we can provide thermodynamic explanations based on the behaviour of macro-level quantities like temperature and energy describing these systems within the domain of equilibrium statistical mechanics. While non-equilibrium and equilibrium statistical mechanics and phenomenological thermodynamics are distinct theories, minimal thermodynamics is a coarse-grained level of description within the framework of equilibrium statistical mechanics. It qualifies as ‘thermodynamics’ especially in virtue of its use of macroscopic coarse-grained quantities partially autonomous from the microscopic variables characterising purely statistical descriptions. We maintain that the picture we develop is the best way to make sense of the notion of gravitational thermodynamics.
@@
Second, analysing SGS allows us to draw general lessons about thermodynamics and statistical mechanics and to study the impact of idealisations in astrophysical models. The case study supports a more liberal approach to concepts such as equilibrium (Callender, 2001), brings out considerations on the role of unconventional properties like negative heat capacity on thermodynamics and statistical physics, and prompts the development of a novel minimal framework for thermodynamics that accounts for thermodynamic descriptions in between the purely statistical and phenomenological thermodynamics level. These results have a double outcome: (a) we develop a useful notion of thermodynamics beyond phenomenological thermodynamics and control theory (Wallace, 2014) with further possible applications; (b) we improve our understanding of equilibrium statistical mechanics, as we show how we can effectively apply equilibrium statistical mechanics to these unconventional systems if only we take a less stringent approach to features like equilibrium and stability. In addition, analysing SGS brings interesting considerations about the role of idealisations and what can we learn from idealised models via de-idealisation methods.
@@ 2 Atypical Features of Self-gravitating Systems
Self-gravitating systems (SGS) are systems bound by their own gravity. Globular clusters are the prime examples of these systems. They are spheroidal conglomerations of self-gravitating stars, composed of tens of thousands to millions of elements. The physics of these systems can be in theory described to a high approximation via Newtonian gravitation. However, this million-body problem, as Heggie and Hut (2003) call it, is obviously practically intractable in that way. Because of this, statistical physics methods are employed to model these systems. [^5] For a classic reference on galactic dynamics see Binney and Tremaine (2011). SGS can be modelled like gases in which stars take the part of gas molecules, but this is not a perfect analogy. The crucial difference between SGS and conventional terrestrial systems to which statistical physics and thermodynamics are applied stems from the nature of the gravitational potential, which is long-range, and from the fact that gravity is universally attractive and does not possess any natural screening-off mechanism. As a consequence, unlike in the case of conventional gas molecules in which only short-range forces are dominant, the potential acting on any given constituent of SGS largely originates from the components that are not in its vicinity. These unconventional conditions determine several well-known kinds of problematic behaviour both from the point of view of statistical physics and thermodynamics: [^6] See Lynden-Bell and Lynden-Bell (1977), Padmanabhan (1990), Binney and Tremaine (2011), Swendsen (2012, §13), Heggie and Hut (2003). For a short overview see Hut (1997).
@@
    Non-extensivity: Due to the long-range nature of gravity, functions such as energy and entropy are not extensive (in the statistical description), i.e. don’t depend linearly on the size of the system, unlike for conventional systems. As reported by Callender (2011), many physics textbooks take extensivity as an axiomatic feature of statistical mechanics and thermodynamics. Even without taking it as axiomatic, it can be noticed that without extensivity we cannot derive what are regarded as key thermodynamic relations within statistical mechanics such as $dU=TdS-PdV+\mu dN$.
@@ 3.2 Callender on self-gravitating systems
Moving to ensemble inequivalence, recall that, unlike in conventional thermodynamics, the ensembles turn out to be generally inequivalent within SGS. This requires us to modify statistical mechanics within SGS, and in general, the equivalence of ensembles should not be taken anymore as a general principle. Callender however suggests that this is not really problematic and does not threaten the application of thermodynamics. Equivalence is a surprising feature of non-long-range interacting systems that emerges in the thermodynamic limit $N,V\rightarrow\infty$ (constant $N/V$), while there is no reason to think it is an essential prerequisite for statistical mechanics or thermodynamics.
@@
Concerning feature (3), i.e. non-extensivity, Callender points out that if we consider the mean-field approximation with limit $N\rightarrow\infty$ (at volume fixed) we regain extensivity. In that limit, energy and entropy become extensive (proportional to N) again. [^16] The limit is applied to the Hamiltonian $H=\sum_{i=1}^{N}\frac{\boldsymbol{p}_{i}^{2}}{2m}+\frac{1}{2N}\sum_{i,j=1}^{N}V(\boldsymbol{q}_{i}-\boldsymbol{q}_{j})$. An important point is that the limit employed is not the thermodynamic limit defined earlier. We mentioned how the absence of this limit within SGS was stressed by Robertson as one of the main reasons why the inapplicability of thermodynamics to SGS is unsurprising. Related to this point, Callender claims that the thermodynamic limit $N,V\rightarrow\infty$ (constant $N/V$) is neither necessary nor sufficient for thermodynamic behaviour. [^17] See also Butterfield (2011) on limits not being necessary for emergence. Most importantly, he suggests that it might be wrong to think that one kind of limit is the right one generally speaking, and there is no a priori reason to prefer that specific limit in every context. Instead, he suggests that we could take the limit $N\rightarrow\infty$ as the most suitable for SGS to recover thermodynamics since that limit allows us to regain extensivity. For this reason, we can call $N\rightarrow\infty$ the ‘gravitational thermodynamic limit’. [^18] On the use of the mean-field and $N\rightarrow\infty$ to regain extensivity and the ability to define properly thermodynamic quantities, similar to what happens in short-range systems in the $N,V\rightarrow\infty$ (constant $N/V$) thermodynamic limit, see Padmanabhan (1990) and Katz (2003). On the thermodynamic limit $N\rightarrow\infty$ for SGS see also Chavanis (2006, §7) and Chavanis (2013, Appendix A). On the use of alternative limits see De Vega and Sanchez (2002, 2006) who employ the dilute limit $N\rightarrow\infty$, $V\rightarrow\infty$, $N/V^{1/3}=fixed$ in ${}^{,}$ [^19] However, Callender is worried that this move could be begging the question since we are deeming the $N\rightarrow\infty$ as the right thermodynamic limit in this context by starting from the assumption that extensivity must hold. We return to this point and refine this limit in §4.
@@
Concerning features (4) and (5), i.e. negative heat capacity and the strictly related instability and lack of equilibrium, the situation is more complex. Are there equilibrium states for SGS? Callender briefly considers the states mentioned before: a Dirac peak in which all the particles sit at the same material point when the (confined) system is attached to a heat bath, or an endlessly unbound shrinking-core system when the system is isolated. Should these exotic and arguably unphysical states count as equilibrium states in any way? Callender does not take a side and instead chooses another route. He suggests that there is no equilibrium for SGS only if we adopt a strict notion of equilibrium. That is, he stresses that equilibrium is an idealisation and a matter of scales: “Equilibrium, in thermodynamics and statistical mechanics, is an idealization. […] there can’t really be any truly stationary states. […] Equilibrium holds only with respect to certain observables, spatial scales and temporal scales” (Callender, 2011, p. 967). Hence, considering the scales involved in SGS, he suggests taking the metastable states of the systems as a viable basis for equilibrium.
@@ 4.1 Statistical physics of idealised self-gravitating systems
Let’s start by stating our position own position in the context of the Callender-Roberson debate. We agree with part of Callender’s conclusions, in particular we concur that both introducing short-distance cutoffs to avoid divergence of the partition function and the phenomenon of ensemble inequivalence are unproblematic steps. We also adopt for now the proposal of introducing an ideal box to model these systems to avoid density of state divergence, although the next subsection shows that things are more complex. We agree with the claims that extensivity can be regained in the appropriate but unconventional limit and that metastable states provide a useful basis for equilibrium within SGS, but we elaborate further on both topics. We also clarify the role of stability and negative heat capacity, showing that these systems are effectively stable and in equilibrium at the right scale.
The absence of equilibrium and stability are central to Robertson’s argument, hence these are our starting points. As maintained by Callender (2001, 2011) equilibrium is generally scale relative, and perfect equilibrium is never reached even in conventional systems. Temporary fluctuations away from statistical equilibrium are to be expected and thus real systems are always in equilibrium relative to a certain scale. [^20] This point aligns well with Ladyman’s and Ross’s thesis of the scale relativity of ontology: “what (really, mind-independently) exists should be relativized to (real, mind-independent) scales at which nature is measurable.” (Ladyman and Ross, 2007, p. 200). If ontology is scale relative, then what it is to be equilibrium is to be equilibrium at a certain scale. On scale relativity see also Franklin and Robertson (2021) and Ladyman and Lorenzetti (2027). Drawing on this thesis, our central points are that (i) equilibrium states can be defined for SGS in the form of local equilibrium if considered in the appropriate regime and limit, as local entropy maxima instead of global maxima, and (ii) these equilibria are metastable states which are effectively stable relative to the right scale, as core collapse is triggered only in certain parameter regimes. While metastability refers to the stability of these states under minor perturbations, we call these entropy maxima ‘quasi-equilibrium states’ to stress that they represent local equilibria that can nevertheless act as proper equilibria over the appropriate time scale, especially as SGS metastable states have extremely long lifetimes. [^21] On ‘quasi-equilibrium states’ we follow the terminology by Heggie and Hut (2003).
@@
Note that they call this feature ‘thermodynamic’ equilibrium but what they are really employing is statistical mechanical equilibrium, as per the definitions in §2. Indeed the same equilibrium is elsewhere referred to as statistical equilibrium (Chavanis, 2006, §4.3), but most importantly this is the kind of equilibrium calculated e.g. from the phase volume and the distribution function. [^24] Physicists working on the subject often call this thermodynamic equilibrium because they seem to identify thermodynamics with equilibrium statistical mechanics to a certain extent and because their focus is on studying thermodynamic potentials in equilibrium. Their rationale for labelling this ‘thermodynamic’ is vindicated within our minimal approach to thermodynamics introduced in §5.
@@
Figure 1: Series of equilibria $(E,\beta)$ for classical isothermal spheres (a thermodynamic-kind diagram). (Chavanis et al., 2002)
Figure 1: Series of equilibria $(E,\beta)$ for classical isothermal spheres (a thermodynamic-kind diagram). (Chavanis et al., 2002)
Figure 2: Basin of attraction in canonical ensemble. Isothermal spheres after the first turning point of the spiral are unstable in this ensemble. Depending on their position on the spiral and the initial perturbation, they can either relax towards the local maximum of free energy with same temperature ($\bullet$) or undergo a gravitational collapse ($\triangle$). (Chavanis et al., 2002)
This behaviour can be described at different levels of description, which also corresponds to different levels of precision and complexity. First, we can describe it within kinetic theory, where the series of equilibria can be calculated with some difficulty via the Fokker-Planck equation for the one-particle distribution function $f(\textbf{x},\textbf{p},t)$ (Padmanabhan (1990, §4), Chavanis (2006, §4.4, §7)): [^26] See also Chavanis (2013).
@@
Second, slightly more easily, we can calculate the variation of entropy within equilibrium statistical mechanics in the canonical and microcanonical ensemble, respectively from the density of states $g(E)$ and the partition function $Z(\beta)$ (Horwitz and Katz (1977), Padmanabhan (1990)). [^27] More precisely, for instance, Katz (2003, §2) employs a steepest descent technique starting from the density of states function. See also Chavanis (2002) and De Vega and Sanchez (2002). Finally, at an even simpler and coarser-grained level, we can maximize the thermodynamic potential (Chavanis, 2006, §2-4) and employ Poincaré linear series of equilibria as developed by Katz (1978, 1980, 2003). [^28] See Heggie and Hut (2003, §17) and Chavanis (2006, §4.7) for an introduction. This method then allows us to analyse the stability of SGS by focusing on the macro-quantities involved in our diagrams and on the topology of the curve $\beta/E$. Systems described by those parameters are stable along the curve until they reach a bifurcation or turning point, which is represented in the diagrams by vertical lines tangent to the equilibrium states line. We can use this method to study instabilities and predict in which parameter regimes they occur, matching the results of kinetic theory and equilibrium statistical mechanics based on $g(E)$ and $Z(\beta)$ but wholly relying on macro-level quantities defined over (scale-relative) equilibrium states within the theoretical framework of equilibrium statistical mechanics (Chavanis, 2006, §7).292929 To anticipate, Katz’s theory will provide in Section 5 the basis for the formulation of a minimal notion of thermodynamics that applies to SGS over merely statistical descriptions.
@@
This presentation shows that equilibrium and stability hold within SGS and that equilibrium statistical mechanics plays a crucial role in the physics of SGS. It is also important to stress the long lifetime of metastable states in SGS: [^30] See also Katz (2003, §2.6).
@@
The lifetime of metastable states (local entropy maxima) scales as $exp(N)$ due to the long-range nature of the interaction. Therefore, the importance of these metastable states is considerable and they cannot be simply ignored. Metastable states are in fact stable and they correspond to observed structures in the universe such as globular clusters. (Chavanis, 2005, pp. 135)
@@
The general conclusion to draw is that metastable states provide stable equilibrium states in SGS within certain parameter regimes and time scales. These metastable states have long lifetimes compared to the evolution of these systems, a feature which can be related to the consideration that equilibrium is always relative to a certain time scale. Hence these technically quasi-equilibrium states can be regarded as effectively equilibrium states at the appropriate time scale. Collapses and instabilities take place and play an important role in the dynamics of SGS, but for the given reasons we can argue that equilibrium and stability can both be recovered within SGS in certain domains. It is worth noting that two distinct scales are operating in this context: a time scale (quasi-equilibrium) and a perturbation scale (metastability). These conclusions contrast Robertson’s claims that equilibrium is not attainable within SGS and that only non-equilibrium statistical mechanics is employed and support instead the thesis that equilibrium statistical mechanics is applicable in this context, although this is not the full picture yet.
@@
Having analysed equilibrium and stability, we now point out another key aspect related to the applicability of equilibrium statistical mechanics to SGS. As reported in §3.1 Robertson argues that “the thermal physics of SGS never abstracts away from to macroscopic bulk variables from the microvariables – i.e. the position and momenta of the individual stars – and probability distributions over these microvariables” (Robertson, 2019, p. 1795). In contrast, we highlight how macroscopic quantities actually play a role in the physical picture we have presented, especially variables such as $E$, $N$, and $V$ that are used to define the density of states $\Omega(E,N,V)$ and entropy $S(E,N,V)$. These quantities are central to equilibrium statistical mechanics operations. But the list does not stop here. As mentioned, Callender (2011) rightly stresses the ability of the mean-field approximation with $N\rightarrow+\infty$ to regain us extensivity of energy. [^31] More precisely, the limit is usually taken with fixed $\eta=\beta GM/R$ and $\Lambda=-ER/GM^{2}$ (Chavanis, 2006, p. 3127). However, employing this approximation and limit brings more than that. In this situation, we can define several macroscopic quantities which are characteristic of the physics of equilibrium statistical mechanics, as Katz (2003, §3.1) remarks: “Besides E, N, V and $S(E,N,V)$ there are other thermodynamic functions that give a physical content to the results […] in particular the derivatives of S with respect to E, N, V” such as $\partial S/\partial E=\beta$, $\partial S/\partial N=-\alpha$, $\partial S/\partial V=\beta P_{b}$ (where $P_{b}$ is pressure on the boundary), from which follows:

### v2

@@ Abstract
The use of statistical methods to model gravitational systems is crucial to physics practice, but the extent to which thermodynamics and statistical mechanics genuinely apply to these systems is a contentious issue. This paper provides new conceptual foundations for gravitational thermodynamics by reconsidering the nature of key concepts like equilibrium and advancing a novel way of understanding thermodynamics. The challenges arise from the peculiar characteristics of the gravitational potential, leading to non-extensive energy and entropy, negative heat capacity, and a lack of standard equilibrium. Hence it has been claimed that only non-equilibrium statistical mechanics is warranted in this domain, whereas thermodynamics is inapplicable. We argue instead that equilibrium statistical mechanics applies to self-gravitating systems at the relevant scale, as they display equilibrium in the form of metastable quasi-equilibrium states. We then develop a minimal framework for thermodynamics that can be applied to these systems and beyond. Thermodynamics applies in the sense that we can devise macroscopic descriptions and explanations of the behaviour of these systems in terms of coarse-grained quantities like energy and temperature within equilibrium statistical mechanics.
@@ 1 Introduction
Second, although full-blown phenomenological thermodynamic descriptions are unsuitable in this domain, we develop what we call a ‘minimal framework for thermodynamics’ and show how a notion of thermodynamics applies to SGS. In fact, we can provide thermodynamic explanations based on the behaviour of macro-level quantities like temperature and energy describing these systems within the domain of equilibrium statistical mechanics. While non-equilibrium and equilibrium statistical mechanics and phenomenological thermodynamics are distinct theories, minimal thermodynamics is a coarse-grained level of description within the framework of equilibrium statistical mechanics. It qualifies as ‘thermodynamics’ especially in virtue of its use of macroscopic coarse-grained quantities partially autonomous from the microscopic variables characterising purely statistical descriptions. We maintain that the picture we develop is the best way to make sense of the notion of gravitational thermodynamics.
@@
Second, analysing SGS allows us to draw general lessons about thermodynamics and statistical mechanics and to study the impact of idealisations in astrophysical models. The case study supports a more liberal approach to concepts such as equilibrium (Callender, 2001), brings out considerations on the role of unconventional properties like negative heat capacity on thermodynamics and statistical physics, and prompts the development of a novel minimal framework for thermodynamics that accounts for thermodynamic descriptions in between the purely statistical and phenomenological thermodynamics level. These results have a double outcome: (a) we develop a useful notion of thermodynamics beyond purely phenomenological thermodynamics with further possible applications; (b) we improve our understanding of equilibrium statistical mechanics, as we show how we can effectively apply equilibrium statistical mechanics to these unconventional systems if only we take a less stringent approach to features like equilibrium and stability. In addition, analysing SGS brings interesting considerations about the role of idealisations and what can we learn from idealised models via de-idealisation methods.
@@ 2 Atypical Features of Self-gravitating Systems
Self-gravitating systems (SGS) are systems bound by their own gravity. Globular clusters are the prime examples of these systems. They are spheroidal conglomerations of self-gravitating stars, composed of tens of thousands to millions of elements. The physics of these systems can be in theory described to a high approximation via Newtonian gravitation. However, this million-body problem, as Heggie and Hut (2003) call it, is obviously practically intractable in that way. Because of this, statistical physics methods are employed to model these systems. [^5] For a classic reference on galactic dynamics see Binney and Tremaine (2011). SGS can be modelled like gases in which stars take the part of gas molecules, but this is not a perfect analogy. The crucial difference between SGS and conventional terrestrial systems to which statistical physics and thermodynamics are applied stems from the nature of the gravitational potential, which is long-range, and from the fact that gravity is universally attractive and does not possess any natural screening-off mechanism. As a consequence, unlike in the case of conventional gas molecules in which only short-range forces are dominant, the potential acting on any given constituent of SGS largely originates from the components that are not in its vicinity. These unconventional conditions determine several well-known kinds of problematic behaviour both from the point of view of statistical physics and thermodynamics: [^6] In presenting these points I follow the exposition in Callender (2011) who introduced these features in the philosophical literature. These points can be found widely in the physics literature, see Lynden-Bell and Lynden-Bell (1977), Padmanabhan (1990), Binney and Tremaine (2011), Swendsen (2012, §13), Heggie and Hut (2003). For an overview see Hut (1997).
@@
    Non-extensivity: Due to the long-range nature of gravity, functions such as energy and entropy are not extensive (in the statistical description), i.e. don’t depend linearly on the size of the system, unlike for conventional systems. As reported by Callender (2011), many physics textbooks take extensivity as an axiomatic feature of statistical mechanics and thermodynamics. Even without taking it as axiomatic, it can be noticed that without extensivity we cannot derive what are regarded as key thermodynamic relations within statistical mechanics such as $U=TS-PV+\mu N$.
@@ 3.2 Callender on self-gravitating systems
Moving to ensemble inequivalence, recall that, unlike in conventional thermodynamics, the ensembles turn out to be generally inequivalent within SGS. This requires us to modify statistical mechanics within SGS, and, in general, the equivalence of ensembles can no longer be considered a general principle. Callender however suggests that this is not problematic and does not threaten the application of thermodynamics. He argues that equivalence is a surprising feature of non-long-range interacting systems that emerges in the thermodynamic limit $N,V\rightarrow\infty$ (constant $N/V$), while there is no reason to think it is an essential prerequisite for statistical mechanics or thermodynamics. [^16] A point we are sympathetic to.
@@
Concerning feature (3), i.e. non-extensivity, Callender points out that we can regain extensivity if we consider mean-field models and follow the so-called ‘Kac’ prescription, appropriately rescaling the Hamiltonian for self-gravitating systems and applying the $N\rightarrow\infty$ limit (at fixed volume). [^17] The limit is applied to the Hamiltonian $H=\sum_{i=1}^{N}\frac{\boldsymbol{p}_{i}^{2}}{2m}+\frac{1}{2N}\sum_{i,j=1}^{N}V(\boldsymbol{q}_{i}-\boldsymbol{q}_{j})$. Note that only extensivity is regained whereas energy is still non-additive. See also Campa et al. (2014, p. 32) on the Kac prescription. In that limit, energy and entropy become extensive (proportional to N) again. [^18] In §4.1 we present yet another sense in which the $N\rightarrow\infty$ limit matters to SGS physics. An alternative limit is also introduced with a similar purpose by De Vega and Sanchez (2002, 2006) who employ the dilute limit $N\rightarrow\infty$, $V\rightarrow\infty$, $N/V^{1/3}=fixed$ in astrophysics. An important point is that these limits are not the thermodynamic limit $N,V\rightarrow\infty$ (constant $N/V$) defined earlier. We mentioned how Robertson stressed the absence of this limit within SGS as one of the main reasons why the inapplicability of thermodynamics to SGS is unsurprising. Related to this point, Callender claims that the thermodynamic limit $N,V\rightarrow\infty$ (constant $N/V$) is neither necessary nor sufficient for thermodynamic behaviour. [^19] See also Butterfield (2011) on limits not being necessary for emergence. Most importantly, he suggests that it might be wrong to think that one kind of limit is the right one generally speaking, and there is no a priori reason to prefer that specific limit in every context. Instead, he suggests that we could take the limit $N\rightarrow\infty$ as the most suitable for SGS to recover thermodynamics since that limit allows us to regain extensivity. For this reason, we can call $N\rightarrow\infty$ the ‘gravitational thermodynamic limit’. [^20] However, Callender is worried that this move could beg the question since we are deeming the $N\rightarrow\infty$ as the right thermodynamic limit in this context by starting from the assumption that extensivity must hold. We return to this point and refine this limit in §4.
@@
Concerning features (4) and (5), i.e. negative heat capacity and the strictly related instability and lack of equilibrium, the situation is more complex. Are there equilibrium states for SGS? Callender briefly considers the states mentioned before: a Dirac peak in which all the particles sit at the same material point when the (confined) system is attached to a heat bath, or an endlessly unbound shrinking-core system when the system is isolated. Should these exotic and arguably unphysical states count as equilibrium states in any way? Callender does not take a side and instead chooses another route. He suggests there is no equilibrium for SGS only if we adopt a strict notion of equilibrium. That is, he stresses that equilibrium is an idealisation and a matter of scales: “Equilibrium, in thermodynamics and statistical mechanics, is an idealization. […] there can’t really be any truly stationary states. […] Equilibrium holds only with respect to certain observables, spatial scales and temporal scales” (Callender, 2011, p. 967). Hence, considering the scales involved in SGS, he suggests taking the metastable states of the systems as a viable basis for equilibrium.
@@ 4.1 Statistical physics of idealised self-gravitating systems
Let’s start by stating our position in the context of the Callender-Robertson debate. We agree with part of Callender’s conclusions, in particular we concur that both introducing short-distance cutoffs to avoid divergence of the partition function and the phenomenon of ensemble inequivalence are unproblematic steps. We also adopt for now the proposal of introducing an ideal box to model these systems to avoid density of state divergence, although the next subsection shows that things are more complex. We agree with the claims that extensivity can be regained in the appropriate but unconventional limit and that metastable states provide a useful basis for equilibrium within SGS, but we elaborate further on both topics. We also clarify the role of stability and negative heat capacity, showing that these systems are effectively stable and in equilibrium at the right temporal and spatial scale.
The absence of equilibrium and stability are central to Robertson’s argument, hence these are our starting points. As maintained by Callender (2001, 2011) equilibrium is generally scale relative, and perfect equilibrium is never reached even in conventional systems. Temporary fluctuations away from statistical equilibrium are to be expected and thus real systems are always in equilibrium relative to a certain scale. Drawing on this thesis, our central points are that (i) equilibrium states can be defined for SGS in the form of local equilibrium if considered in the appropriate regime and limit, as local entropy maxima instead of global maxima, and (ii) these equilibria are metastable states which are effectively stable relative to the right scale, as core collapse is triggered only in certain parameter regimes. While metastability refers to the stability of these states under minor perturbations, we call these entropy maxima ‘quasi-equilibrium states’ to stress that they represent local equilibria that can nevertheless act as proper equilibria over the appropriate time scale, especially as SGS metastable states have extremely long lifetimes.
@@
Note that they call this feature ‘thermodynamic’ equilibrium but what they are really employing is statistical mechanical equilibrium, as per the definitions in §2. Indeed the same equilibrium is elsewhere referred to as statistical equilibrium (Chavanis, 2006b, §4.3), but most importantly this is the kind of equilibrium calculated e.g. from the phase volume and the distribution function. [^23] Physicists working on the subject often call this thermodynamic equilibrium because they seem to identify thermodynamics with equilibrium statistical mechanics to a certain extent and because their focus is on studying thermodynamic potentials in equilibrium. Their rationale for labelling this ‘thermodynamic’ is vindicated within our minimal approach to thermodynamics introduced in §5.
@@
Figure 1: Series of equilibria $(E,\beta)$ for classical isothermal spheres (a thermodynamic-kind diagram). (Chavanis et al., 2002, Figure 1)
Figure 2: Basin of attraction in canonical ensemble. Isothermal spheres after the first turning point of the spiral are unstable in this ensemble. Depending on their position on the spiral and the initial perturbation, they can either relax towards the local maximum of free energy with same temperature ($\bullet$) or undergo a gravitational collapse ($\triangle$). (Chavanis et al., 2002, Figure 16)
This behaviour can be described at different levels of description, which also corresponds to different levels of precision and complexity. First, we can describe it within kinetic theory, where the series of equilibria can be calculated with some difficulty via the Fokker-Planck equation for the one-particle distribution function $f(\textbf{x},\textbf{p},t)$ (Padmanabhan (1990, §4), Chavanis (2006b, §4.4, §7)): [^25] See also Chavanis (2013).
@@
Second, slightly more easily, we can calculate the variation of entropy within equilibrium statistical mechanics in the canonical and microcanonical ensemble, respectively from the density of states $g(E)$ and the partition function $Z(\beta)$ (Horwitz and Katz (1977), Padmanabhan (1990)). [^26] More precisely, for instance, Katz (2003, §2) employs a steepest descent technique starting from the density of states function. See also Chavanis (2002) and De Vega and Sanchez (2002). Finally, at an even simpler and coarser-grained level, we can maximize the thermodynamic potential (Chavanis, 2006b, §2-4) and employ Poincaré linear series of equilibria as developed by Katz (1978, 1980, 2003). [^27] See Heggie and Hut (2003, §17) and Chavanis (2006b, §4.7) for an introduction. This method then allows us to analyse the stability of SGS by focusing on the macro-quantities involved in our diagrams and on the topology of the curve $\beta/E$. Systems described by those parameters are stable along the curve until they reach a bifurcation or turning point, which is represented in the diagrams by vertical lines tangent to the equilibrium states line. We can use this method to study instabilities and predict in which parameter regimes they occur, matching the results of kinetic theory and equilibrium statistical mechanics based on $g(E)$ and $Z(\beta)$ but wholly relying on macro-level quantities defined over (scale-relative) equilibrium states within the theoretical framework of equilibrium statistical mechanics (Chavanis, 2006b, §7).282828 To anticipate, Katz’s theory will provide in Section 5 the basis for the formulation of a minimal notion of thermodynamics that applies to SGS over merely statistical descriptions.
@@
This presentation shows that equilibrium and stability hold within SGS and that equilibrium statistical mechanics plays a crucial role in the physics of SGS. It is also important to stress the long lifetime of metastable states in SGS: [^29] See also Katz (2003, §2.6). Chavanis (2005) provides a method to estimate the lifetime of metastable states via Kramers formula. See Chavanis (2024) for the detailed calculation of the lifetime of metastable states of a simple kind of SGS (a binary star) close to critical point.
@@
The lifetime of metastable states (local entropy maxima) scales as $exp(N)$ due to the long-range nature of the interaction. Therefore, the importance of these metastable states is considerable and they cannot be simply ignored. Metastable states are in fact stable […]. (Chavanis, 2005, pp. 135)
@@
The conclusion to draw is that metastable states provide stable equilibrium states in SGS within certain parameter regimes and time scales. These metastable states have long lifetimes compared to the evolution of these systems, a feature which can be related to the consideration that equilibrium is always relative to a certain time scale. Hence these technically quasi-equilibrium states can be regarded as effectively equilibrium states at the appropriate time scale. Collapses and instabilities take place and play an important role in the dynamics of SGS, but for the given reasons we can argue that equilibrium and stability can both be recovered within SGS in certain domains. It is worth noting that two distinct scales are operating in this context: a time scale (quasi-equilibrium) and a perturbation scale (metastability). These conclusions contrast Robertson’s claims that equilibrium is not attainable within SGS and that only non-equilibrium statistical mechanics is employed and support instead the thesis that equilibrium statistical mechanics is applicable in this context, although this is not the full picture yet.
Having analysed equilibrium and stability, we now discuss two other key features related to the applicability of equilibrium statistical mechanics to SGS. In particular, we contrast the two interrelated Robertson’s objections that (a) SGS physics does not abstract away from the micro-details of the systems, and that (b) the thermodynamic limit does not exist for SGS.
@@
As reported in §3.1 Robertson argues that “the thermal physics of SGS never abstracts away from to macroscopic bulk variables from the microvariables – i.e. the position and momenta of the individual stars – and probability distributions over these microvariables” (Robertson, 2019, p. 1795). In contrast, we highlight how macroscopic quantities actually play a role in the physical picture we have just presented, especially variables such as $E$, $N$, and $V$ that are used to define the density of states $\Omega(E,N,V)$ and entropy $S(E,N,V)$. These quantities are central to equilibrium statistical mechanics operations.
Furthermore, other macroscopic quantities grounded in equilibrium statistical mechanics can be defined within SGS models. As mentioned, Callender (2011) stresses the ability of the mean-field approximation with $N\rightarrow+\infty$ at fixed volume (via the ‘Kac’ prescription) to regain us extensivity of energy. But the use of such a limit in the mean-field brings more than that. As shown e.g. by Chavanis (2006a, b), by taking the $N\rightarrow+\infty$ limit (with fixed $\eta=\beta GM/R$ and $\Lambda=-ER/GM^{2}$) the mean-field approximation becomes exact, meaning that the total probability distribution for the system factorises, making the N-body distribution function a product of individual mean-field Maxwell-Boltzmann distribution functions. Hence we can apply an equilibrium statistical description, computing entropy maxima using the new distribution. [^30] On the use of ‘thermodynamic’ functions for SGS in the $N\rightarrow+\infty$ limit see also Tatekawa et al. (2005, §6), Casetti and Nardini (2010, §3.1.1 and §3.2.1), Follana and Laliena (2000).
In this context, we can define several macroscopic quantities which are characteristic of the physics of equilibrium statistical mechanics, as Katz (2003, §3.1) remarks: “Besides E, N, V and $S(E,N,V)$ there are other thermodynamic functions that give a physical content to the results […] in particular the derivatives of S with respect to E, N, V” such as $\partial S/\partial E=\beta$, $\partial S/\partial N=-\alpha$, $\partial S/\partial V=\beta P_{b}$ (where $P_{b}$ is pressure on the boundary), from which follows: